OpenStudy (anonymous):

What is the domain of the function f(x) = 2x2 + 7x - 4?

OpenStudy (zzr0ck3r):

is there any reason what so ever that we CANT plug in each and every real number?

OpenStudy (zzr0ck3r):

if f(x) is a polynomial then \(Domain = \mathbb{R}\)

OpenStudy (anonymous):

A. (-,2) B. (-,) C. [4,) D. [-4,) these are the answers to chose from

OpenStudy (zzr0ck3r):

none of those answers make sense

OpenStudy (anonymous):

infinity

OpenStudy (anonymous):

it didnt put it in when i copied it

OpenStudy (zzr0ck3r):

ahh so \[(-\infty, \infty)=\mathbb{R}\]

OpenStudy (zzr0ck3r):

the domain to any polynomial is all real numbers

OpenStudy (anonymous):

i will make a bet my bet is they are asking for the "range" just a guess

OpenStudy (zzr0ck3r):

They always ask this domain question with polynomials at least once. But I would not be surprised if you are right @satellite73

OpenStudy (zzr0ck3r):

well the options for the range and the domain are both given, so ....lol